40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8620     Entry time: Wed May 22 18:24:19 2013     In reply to: 8611
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Violin mode survey 

Quote:

It was too embarassing to see that the actuation frequency was set at the violin mode frequency in order to avoid designing a new filter!?

 Ooops, definitely my bad.  I think I was the one who put in the PRM violin filter, so I should have recognized that frequency.  However, I couldn't think of a reason why violin mode filters should be in the LSC filter banks, since we usually put them in C1:SUS-optic_LSC filter banks. 

Anyhow, so that I don't make a mistake like that again, I was looking through all of the violin mode filters for all the optics, so I could write down the frequencies.  The result: confusion.

Violin filters in C1:SUS-optic_LSC filter banks:

The PRM's violin mode filter is set correctly to 627.75Hz:  elog 8533.

One of BS or SRM has probably been measured (presumably BS), since they have the same filters centered around 645Hz. 

Neither ITM has a violin filter.

The ETMs have violin filters in the 440's, which I assume was correct back in the MOS days, before 2010.

Vio2 filters in C1:SUS-optic_LSC filter banks:

PRM, SRM, BS, ITMX, ITMY:  Centered around 1285 Hz, which matches the violin notch frequencies in the BS and SRM.

ETMY:  Centered around 883.5Hz, which matches the old 440Hz frequency

ETMX:  Centered around 631Hz .  So, this could have been measured, but it was put into the wrong filter module.

 

Koji tells me that we don't really need to worry about all these violin filters unless they are required (as with the PRM and the obnoxious hum a few weeks ago), so I 'm not going to do any measuring / adjusting of these filters for now.

ELOG V3.1.3-